OpenStudy (anonymous):

Find an equation of the tangent line to the given curve at 'a': y=((e^x)/4)-x ; a=0

OpenStudy (anonymous):

take the derivative, replace \(x\) by 0 to find the slope what you wrote is not very clear however is it \[f(x)=\frac{e^x}{4}-x\]?

OpenStudy (anonymous):

Yes, that is correct

OpenStudy (anonymous):

\(f(0)=\frac{1}{4}\) so your point is \((0,\frac{1}{4})\) \[f'(x)=\frac{1}{4}e^x-1\] \[f'(0)\] is your slope, then use the point slope formula

OpenStudy (anonymous):

Thank you. Now, the next question is the same type of question, but the equation is \[y= e ^{x}\] with a=ln3. I'm still trying to find the equation of the tangent line of the curve at 'a'. I know the answer, but I don't know how to get to it.

OpenStudy (anonymous):

dervivative of \(e^x\) is \(e^x\) your point is \((\ln(3), 3)\) since \(e^{\ln(2)}=3\)

OpenStudy (anonymous):

and similarly your slope is \(3\)